OpenStudy (anonymous):

If y varies inversely as x, and y = 3 when x = 6, find the value of x when y = 18

jimthompson5910 (jim_thompson5910):

"If y varies inversely as x" means y = k/x Since "y = 3 when x = 6", we can plug them in to get 3 = k/6 Solve that for k Then use this value of k to find the value of x when y = 18

OpenStudy (anonymous):

whats k ? 18 ?

jimthompson5910 (jim_thompson5910):

yes, k = 18

OpenStudy (anonymous):

3= 18/6

jimthompson5910 (jim_thompson5910):

y = k/x y = 18/x 18 = 18/x 18x = 18 x = 18/18 x = 1 So when y = 18, x = 1

OpenStudy (anonymous):

but these are the answers 36 6 9 1

jimthompson5910 (jim_thompson5910):

Alright, so where's the issue? I wrote the answer above